OpenStudy (anonymous):

The equation of line EF is y = 1 over 2x + 6. Write an equation of a line parallel to line EF in slope-intercept form that contains point (0, −2).

OpenStudy (texaschic101):

y = 1/2x + 6 can you tell me what the slope of this line is ?

OpenStudy (texaschic101):

in y = mx + b form, the slope is in the m position. So, in y = 1/2x + 6, what number is in the m position ?

OpenStudy (anonymous):

idk but these are the answer choices y = −2x − 2 y = negative 1 over 2x + 2 y = 1 over 2x − 2 y = − 2x + 2

OpenStudy (anonymous):

a. y= -2x-2 b. -1/2x+2 c. y=1/2x-2 d. y=-2x+2